# A Neural-Guided Variational Quantum Algorithm for Efficient Sign Structure Learning in Hybrid Architectures

### v1

@@ Abstract
Variational quantum algorithms hold great promise for unlocking the power of near-term quantum processors, yet they are frequently hindered by high measurement costs, barren plateaus, and challenging optimization landscapes. Here, we introduce sVQNHE, a neural-guided variational quantum algorithm that decouples amplitude and sign learning across classical and quantum modules, respectively. Our approach employs shallow quantum circuits composed of commuting diagonal gates to efficiently model quantum phase information, while a classical neural network learns the amplitude distribution and guides circuit optimization in a bidirectional feedback loop. This hybrid quantum-classical synergy not only reduces measurement costs but also achieves high expressivity with limited quantum resources and improves the convergence rate of the variational optimization. We demonstrate the advancements brought by sVQNHE through extensive numerical experiments (including realistic noisy simulations and sampling-based simulations). For the 6-qubit J1-J2 model — a prototypical system with a severe sign problem for Monte Carlo-based methods — it reduces the mean absolute error by $98.9\%$ and suppresses variance by $99.6\%$ relative to a baseline neural network, while requiring nearly $19x$ fewer optimization steps than a standard hardware-efficient VQE. Furthermore, for MaxCut problems on 45-vertex Erdős–Rényi graphs, sVQNHE improves solution quality by $19\%$ and quantum resource efficiency by $85\%$. Importantly, this framework is designed to be scalable and robust against hardware noise and finite-sampling uncertainty, making it well-suited for both current NISQ processors and future high-quality quantum computers. Our results highlight a promising path forward for efficiently tackling complex many-body and combinatorial optimization problems by fully exploiting the synergy between classical and quantum resources in the NISQ era and beyond.
@@ Introduction
Quantum computing promises transformative advances in a variety of applications, from combinatorial optimization to ground-state preparation and dynamical simulation of many-body systems [1, 2]. However, large-scale quantum circuits remain challenging to implement due to hardware limitations, gate errors, and limited qubit coherence [3]. One of the most urgent goals in the noisy intermediate-scale quantum (NISQ) era, and the subsequent early fault-tolerant era [4] (Early FTQC), is to maximize computational utility within these constraints by employing shallow circuits and error-mitigated designs [5].
Among the many strategies proposed to enable practical applications under these conditions, hybrid quantum–classical algorithms, particularly variational quantum algorithms (VQAs) [6, 7, 8], stand out. VQAs combine parameterized quantum circuits (PQC) with classical optimization routines, using quantum circuits as ansatzes to solve problems that can be formulated within a variational framework. Despite initial enthusiasm, standard VQAs face at least three critical challenges. First, optimization challenges, such as barren plateaus [9, 10, 11, 12, 13], where gradients vanish exponentially with circuit depth or system size, severely limit trainability. Second, measurement costs scale poorly with both qubit count and circuit parameters, placing heavy demands on quantum resources. Third, even hardware-efficient circuit designs may require long depths to represent certain wavefunctions, but decoherence makes such computations unreliable in practice [14].
Meanwhile, deep learning has demonstrated remarkable capabilities in learning complex functions and high-dimensional structures from large datasets. In fact, neural networks (NNs) have set new standards for computational efficiency in tasks like quantum simulation [15, 16, 17, 18, 19], which are also commonly associated with quantum computing. A natural question then arises: can quantum resources enhance NN-based variational methods and help push the computational boundary of what is achievable with current and near-future hardware? Previous works have extensively investigated and documented that NNs can efficiently model the wavefunction amplitudes (i.e., essentially the probability distribution), but struggle to capture intricate sign or phase structures in many-body systems [20, 21, 22, 23, 19, 3]. On the other hand, modeling the phase structure of an arbitrary wavefunction can be well approximated with some variations of parameterized Instantaneous Quantum Polynomial (IQP) circuit [24, 25, 26, 27, 28, 29]. This observation motivates one way to combine PQC with NN to jointly model complex wavefunction within a hybrid variational framework. Importantly, such models may exhibit potential quantum advantage if continuous variable IQP (CV-IQP) operations exist for a single iteration [30, 31].
Numerous efforts have explored the combination of quantum computing with deep learning in various settings [32, 33, 34]. An especially relevant example is the variational quantum-neural hybrid eigensolver (VQNHE) [35] algorithm, previously developed by one of us and collaborators. This method uses a neural network to post-process quantum states prepared by a shallow circuit, forming a hybrid quantum-neural ansatz that efficiently expands the accessible quantum state space. However, VQNHE still faces limitations in trainability and incurs non-trivial measurement overhead.
In light of these developments and challenges, we revisit how best to harness both quantum and classical resources within a variational framework, one that maximizes hybrid advantages while mitigating their downsides. To this end, we propose sign-VQNHE (sVQNHE), a hybrid algorithm that comprehensively assuages all the major challenges facing mainstream VQAs: expressiveness, measurement cost, trainability, and robustness (see Table 1). In short, the key innovations of sVQNHE lie in its principled separation of quantum and classical roles: the diagonal layers of the quantum circuit are tasked to model the phase structure (with much reduced measurement costs), while a non-negative neural network only has to learn the amplitude (or effectively, the probability) distribution for improved numerical stability. These components are coupled through an iterative amplitude-transfer mechanism that maps the neural state onto a shallow and classically simulatable quantum layer. Compared to the original VQNHE, this re-design of the hybrid quantum-neural framework not only optimizes expressivity with shallow circuits and substantially improves the measurement costs and the gradient-based convergence rate, but also inherently mitigates sampling noise and decoherence, yielding more robust solutions under realistic conditions (see Results).
During the iterative, bidirectional feedback optimization loop (to train the PQC and NN parameters), if a single diagonal quantum layer cannot fully capture the target wavefunction’s sign structure, we extend the circuit depth by adding a layer of classically simulatable gates along with a new diagonal layer. The classical gate layer approximates the amplitude function learned by the NN by solving a tractable optimization problem, allowing the NN to guide amplitude refinement and assist in layer-wise circuit optimization without encountering barren plateaus. Importantly, because only the parameters of commuting diagonal gates are optimized on the quantum device, the required number of quantum measurements is dramatically reduced compared to the conventional non-commuting circuit ansatzes. In practice, at the last layer of the iterative, bidirectional feedback optimization procedure, one can choose to relax the constraint and allow the neural network to output complex values, thereby enhancing the expressiveness of the hybrid quantum-neural ansatz. Finally, our new algorithm improves the convergence behavior of the original VQNHE by incorporating a stable layer-by-layer training strategy that suppresses numerical instabilities.
This newly proposed iterative and bidirectional feedback loop enables efficient training, expressive state representations, and significant reductions in circuit depth and measurement cost. Our numerical experiments validate these advantages. For instance, in the 6-qubit J1–J2 model, sVQNHE reduces mean absolute error by $98.9\%$ and variance by $99.6\%$ relative to a baseline NN, while achieving nearly 19× faster convergence compared to a hardware-efficient VQE using similar quantum resources. In solving the MaxCut problem on 45-node Erdős–Rényi graphs, sVQNHE improves solution quality by $19\%$ and quantum resource efficiency by $85\%$ over a brickwork ansatz VQE baseline [36].
In summary, sVQNHE offers a scalable and practical path forward for tackling complex many-body quantum simulations and large-scale combinatorial optimization problems by fully leveraging the synergy between classical and quantum computation. As both quantum hardware and neural network models continue to advance, we anticipate that hybrid frameworks like sVQNHE will play an increasingly important role in achieving real-world quantum utility.
Challenge Problem Description sVQNHE Solution
Representation NN-based Variational Monte Carlo struggles to model complex phase structure Delegate the quantum circuit to learn phase structure
Measurement
VQE: $\mathcal{O}(2dn^{k^{\prime}})$, commuting gates compress groups to $\mathcal{O}(n^{k^{\prime}})$ ($k^{\prime}<k$)
VQNHE: $\mathcal{O}(2dn^{k})$, $n$ qubit, $\mathcal{O}(n^{k})$ operators, $d$ parameter
Commuting gates enable simultaneous measurements,
reducing cost to $\mathcal{O}(n^{k})$
Optimization
VQE: barren plateau in deep circuits
VQNHE: potential numerical instability
Neural-guided layer-wise optimization with gradual
amplitude transfer
Noise
VQE: affected by hardware noise and finite sampling noise
VQNHE: NN mitigates errors but amplifies sampling noise
Gradual amplitude transfer reduces sampling noise,
NN mitigates errors
Table 1: Key challenges in existing algorithms and sVQNHE’s mitigation strategies
@@ Results
Figure 1: (1) The algorithm seamlessly integrates quantum circuits with neural networks. The quantum circuits interleaves diagonal layers ($W_{1}\cdots W_{n}$) with shallow, simultable layers ($G_{1}\cdots G_{n}$). Parameters of the $G$ layers are optimized classically using a gradual transfer mechanism, while those of the $W$ layers are adjusted via the parameter shift rule on the actual hardware. Leveraging the commuting properties of the diagonal layers, the number of measurements needed to compute gradients for the $d$ parameters of $W$ drops from $\mathcal{O}(2d)$ to $\mathcal{O}(1)$. (2) Illustration of sVQNHE’s advantages: the star represents the optimal solution, the green circle represents the initial solution space that is easy to explore by the NN, the orange region represents the solution space accessible to the PQC, and the black circle represents the coverage of the hybrid quantum-classical scheme. Among them, (2a) describes the suboptimal situation, that is, PQC cannot connect the classical region with the optimal solution, and the hybrid scheme is difficult to cover the optimal solution; (2b) shows the potential quantum advantage; (2c) describes the transition strategy from (2a) to (2b) through parameter adjustment. (3) Illustration of the energy optimization process using the hierarchical optimization scheme. When entering the next layer, the local optimal solution will be jumped out due to the newly introduced parameters.
We introduce sVQNHE, a hybrid quantum-classical variational algorithm designed to address key limitations of conventional VQAs, including high measurement cost, barren plateaus, limited expressive power with shallow circuits, and noise-induced errors. This framework builds on the original VQNHE approach by introducing a layer-wise, phase-amplitude decoupling strategy[24] (see Fig.1) that enhances trainability, scalability, and sampling robustness.
@@
|$\ket{\psi}=F\prod_{i=0}^{L-1}W_{i}G_{i}\ket{0},$ (1)
where $F$ denotes a post-processing operator defined by a classical NN. The following paragraphs explain the design of the hybrid ansatz and the associated bidirectional optimization strategy. Finally, the algorithm is summarized in Alg. 1.
Quantum Circuit for Phase Structure Learning. The quantum circuit ansatz consists of alternating layers of parameterized gates:$\left(G_{i},W_{i}\right)$. $G_{i}$ are classically simulatable and non-diagonal layers (e.g., $R_{y}$ gates), primarily responsible for shaping the amplitude structure. $G_{1}$ is usually taken to be a layer of Hadamard gates to create a uniform superposition. $W_{i}$ are diagonal gates (e.g., $R_{z}$ and $R_{zz}$) designed to encode the phase structure of the wavefunction. These layers are added iteratively when greater expressivity is needed, and their parameters are optimized in a layer-wise fashion.
Classical Neural Network for Amplitude Modeling. The classical neural network is constrained to produce non-negative outputs, and is used to model the amplitude $F=\sum_{s}f(s)\ket{s}\bra{s}$ | of the quantum state in the computational basis. With the phase information delegated to the quantum circuit, the NN is focused on learning the probability distribution. Prior works have shown that classical neural networks are highly effective at this task, particularly for systems with smooth or structured amplitude landscapes [21, 24, 20].
Bidirectional Feedback Loop: Decoupling Amplitude and Phase. At each iteration $i$, a new block of paramterized quantum layers $(G_{i},W_{i})$ is added to the circuit. The process follows two key steps. (1) Forward Initialization Step: The classical NN operator $F$ from the previous iteration is approximated by the shallow quantum layer $G_{i}$, via a tractable classical optimization procedure as detailed in the Method section. This ensures that $G_{i}$ inherits the amplitude structure previously learned by the NN. (2) Hybrid Optimization: Once $G_{i}$ is initialized, the neural operator $F$ is reset and jointly optimized with the parameters of the new diagonal gate layer $W_{i}$, as done in the original VQNHE algorithm. This hybrid optimization simultaneously refines both the amplitude ($F$) and sign structure ($W_{i}$) in a tightly coupled quantum–classical feedback loop.
@@
6:Define diagonal operator $F_{1}\leftarrow\sum_{s}f(s)\ket{s}\bra{s}$ | $\triangleright$ $f(s)$: NN-predicted amplitude
7:Set $G_{1}\leftarrow H^{\otimes n}$triangleright Uniform superposition
@@
[…]
@@
12:          Update $\theta^{g_{l}}$ to minimize $\|G_{l}-F_{l-1}\|$triangleright Gradual Transfer Mechanism, see Method for detail
@@
14:       $E_{l}\leftarrow 0$triangleright Reset layer energy
@@
17:          Apply operator: |$\ket{\psi_{l}}\leftarrow W_{l}(\theta^{w_{l}})G_{l}(\theta^{g_{l}})\ket{\psi_{l-1}}$
18:          Construct hybrid state: |$\ket{\psi_{l,f}}\leftarrow F_{l}\ket{\psi_{l}}$
@@
24:          Update parameters: $\theta^{w_{l}}$triangleright Measurement Strategy, see Method for detail
25:       until $|\Delta E_{l}|<\epsilon_{\text{conv}}$triangleright Energy convergence
@@
29:return $E_{L}$triangleright Final optimized energy
Adaptive Hybrid Ansatz with Controlled Expressivity
Conventional variational algorithms, such as the VQE and the QAOA, typically rely on predefined circuit ansatzes [37, 38]. These ansatzes, often inspired by physical principles like Trotterized dynamics in QAOA, impose a static architecture that can limit their effectiveness. For example, if the ansatz fails to capture the complex entanglement or sign structure of the target ground state, the optimization may converge to suboptimal local minima. To address this, adaptive methods like adapt-VQE [39, 10] iteratively construct the ansatz by selecting gates from a predefined pool. However, these approaches encounter significant hurdles: the gate pool may not include operations capable of generating sufficient entanglement, thus constraining expressiveness, and the iterative gate selection process requires a large number of quantum measurements, making it impractical for large-scale systems.
In contrast, VQNHE and sVQNHE employ a hybrid quantum-classical ansatz that effectively leverages both quantum and classical resources to represent a quantum state. This approach significantly expands the accessible quantum states with shallow quantum circuits. Nevertheless, the original VQNHE still assumes a fixed ansatz approach, which could clearly be optimized to further reduce circuit depth. To this end, sVQNHE adopts a neural-guided framework to systematically extend the circuit’s depth along with controlled expressiveness. Its shallow quantum layers $G_{l}$ are sequentially determined under the gradual guidance of a classical neural network.
A key feature of sVQNHE is its ability to achieve high expressiveness and stable convergence even with simple layers of $G_{i}$. For instance, take $G_{i}$ to be a layer of $R_{y}$ rotations on each qubit, can already enhannce the ansatz’s flexibility. This capability is rigorously supported by Theorem 1 (Expressiveness of the Quantum Circuit Ansatz), which analyzes the Lie algebra generated by specific sets of operators [40, 41, 12, 42, 43]. Consider the generator set $\mathcal{G}_{1}=\{Z_{D_{2j}}\mid j=1,2,\ldots,\lfloor\frac{m}{2}\rfloor\}\cup\{Y_{i}\mid i=1,2,\ldots,n\}$, where $Y_{i}$ are Pauli Y operators on individual qubits, and $Z_{D_{2j}}$ are Pauli Z operators on specific qubit subsets. The theorem proves that the corresponding Lie algebra is $\mathfrak{g_{1}}=\mathfrak{su}(2^{n-1})\oplus\mathfrak{su}(2^{n-1})$, a highly expressive structure with a dimension far exceeding that of alternative generator sets (e.g., $\mathcal{G}_{2}$). This large Lie algebra indicates that even layers primarily composed of $R_{y}$ gates, can explore a vast state space, enhancing the ansatz’s ability to approximate complex quantum states efficiently.
Moreover, the design of the quantum layers $G_{l}$ in sVQNHE is guided by structural constraints that ensure computational efficiency and scalability. Proposition 1 (Commuting Block Structure) provides a critical insight: if a block diagonal matrix $M=M_{1,1}\oplus\cdots\oplus M_{m,m}$ has distinct eigenvalues across its blocks, any matrix commuting with $M$ must preserve the same block structure. In the context of sVQNHE, this proposition informs the construction of $G_{l}$ layers. By respecting this commuting block structure, sVQNHE can increase circuit depth in a controlled manner, isolating commuting blocks (e.g., $W_{i}$).
Unlike rigid ansatzes (e.g., QAOA’s Trotterized structure) or gate-pool-dependent methods (e.g., adapt-VQE), sVQNHE avoids both architectural rigidity and measurement bottlenecks, achieves higher expressivity with less quantum resource demands.
Efficient Measurement and Trainability via Diagonal Gates and Neural Guidance
@@
The sVQNHE framework addresses these limitations through two synergistic innovations. First, it employs commuting diagonal gates, such as $R_{z}$ and $R_{zz}$ rotations, which enable simultaneous measurement of expectation values in one basis. This compresses the per-operator measurement cost to $\mathcal{O}(1)$, reducing the total overhead to $\mathcal{O}(\text{poly}(n))$ for $\text{poly}(n)$ operators (see details in the method section). This stands in stark contrast to the multiplicative $\mathcal{O}(2d\cdot\text{poly}(n))$ scaling of traditional VQNHE and conventional VQAs, allowing sVQNHE to tackle large-scale problems efficiently.
Secondly, our framework incorporates layer-wise neural guidance to decompose the optimization process into manageable subproblems, addressing a critical limitation of traditional layer-wise VQAs[44]. In conventional layer-wise VQAs, the absence of a global perspective on the optimization landscape renders sequential layer optimization prone to converging at local minima. In contrast, the sVQNHE framework employs a classical neural network to guide the optimization of each quantum layer, producing a hybrid state $|\psi_{l,f}\rangle$ that is more expressive than the state $|\psi_{l}\rangle$ obtained from hierarchical quantum circuits alone. This neural guidance not only helps avoid suboptimal local minima but also focuses parameter updates on one specific block each time, such as $W_{l}$ at the $l$-th iteration. This approach ensures that gradients remain measurable even in deep circuits, since the effective parameter space at each optimization step can always be restricted to a minuscule fraction of the full deep circuit. Together, these strategies circumvent the dual bottlenecks of measurement overhead and barren plateaus, establishing a scalable pathway for variational quantum algorithms in practical applications.
@@
$\text{Var}(\hat{E}_{P_{i}})=\text{Var}\left(\bra{\psi_{L,f}}P_{i}\ket{\psi_{L,f}}\right)\cdot\text{Var}(g^{-1}(s,L)),$ (2)

### v2

@@ Abstract
Variational quantum algorithms are a leading candidate for near-term quantum computing, yet their scalability is hindered by high measurement costs, barren plateaus, and noise sensitivity. These challenges stem largely from requiring quantum circuits to learn both probability amplitudes and phase (sign) structure simultaneously—a demanding regime that increases circuit depth and sampling overhead. Here we introduce sVQNHE, a hybrid quantum–classical framework that explicitly separates these roles: a classical neural network models the amplitude distribution, while shallow, commuting diagonal quantum circuits learn only the phase structure. Furthermore, sVQNHE employs an iterative amplitude-transfer mechanism that progressively integrates the classical distribution into the quantum ansatz—while leveraging commuting diagonal gates to reduce gradient measurement cost to $\mathcal{O}(1)$ per operator—ensuring the neural network models only a residual correction on a highly trainable energy landscape with polynomially bounded gradient variance. Across frustrated quantum spin models, molecular Hamiltonians, and combinatorial optimization benchmarks, sVQNHE significantly outperforms conventional variational quantum eigensolvers with comparable parameter counts, converging faster and demonstrating improved robustness under finite sampling and noise. Notably, using qubit-efficient encodings, sVQNHE matches the state-of-the-art classical Free-Energy Machine on a 1,485-vertex MaxCut instance using Pauli Correlation Encoding, and substantially outperforms all baselines on Maximum Clique problems with 135 vertices. These results establish that explicit amplitude–phase separation, with quantum resources dedicated mainly to learning phase structure, provides a principled and scalable design strategy for hybrid quantum–classical algorithms under near-term noisy hardware constraints.
@@ Introduction
Hybrid quantum–classical approaches represent a promising strategy for extracting computational value from noisy intermediate-scale quantum (NISQ) and the subsequent early fault-tolerant quantum computing (Early FTQC) devices[preskill2018quantum, katabarwa2024early, nanda2024explorative]. Central to this paradigm is the idea of variational wavefunction learning, in which a parametrized ansatz is iteratively optimized to minimize a task-specific objective, such as ground-state energy or a combinatorial cost function[peruzzo2014variational, mcardle2020quantum, cerezo2021variational]. Despite substantial progress across physics, optimization, and quantum machine learning, current variational quantum algorithms (VQAs) scale poorly in the presence of noise and sampling constraints[cerezo2021variational, fedorov2022vqe]. In practice, training often becomes unstable as system size grows, the gradient signal rapidly diminishes, and measurement cost increases with circuit complexity[cerezo2021cost, grant2019initialization, holmes2021barren, holmes2022connecting, wang2021noise, ilin2024dissipative].
These failures stem from a profoundly imbalanced division of labor between quantum and classical resources. In most contemporary VQA architectures, from hardware-efficient layers to QAOA-style circuits, the quantum processor is often tasked to optimize a generic ansatz state with both the probability amplitudes and the delicate phase/sign structure to be tuned simultaneously. Capturing realistic amplitude distributions typically demands highly entangling gates and considerable circuit depth[grimsley2019adaptive, chen2020variational], whereas correctly learning the phase is extraordinarily fragile under hardware noise[cerezo2021cost, wang2021noise, ragone2024lie]. As a result, improvements in one aspect often degrade the other, driving the well-documented trade-offs among circuit depth, trainability, sample complexity, and noise sensitivity.
A largely unexplored yet critical question is therefore how to intelligently partition amplitude and phase learning between classical and quantum resources. Suppose the amplitude landscape can be captured classically[bluvstein2021controlling, lange2024architectures, pfau2024accurate, hermann2020deep, carleo2017solving] while reserving quantum resources for the inherently hard phase structure[westerhout2023many, park2025complexity, pagano2020quantum, leontica2024exploring, diez2024connection, szabo2020neural, westerhout2020generalization, nomura2017restricted, cai2018approximating, carleo2017solving, preskill2018quantum]. In that case, the quantum circuit may remain shallow, measurement-efficient, and less affected by noise, without sacrificing overall expressivity. However, realizing such a division of labor poses its own challenges. The wavefunction representation must remain valid, optimization must be stable across heterogeneous parameter spaces, and the classical component must not eliminate the need for quantum computation altogether. Unlike classical variational quantum Monte Carlo methods, where amplitude-phase decomposition is straightforward, integrating quantum circuits introduces unique technical hurdles, such as ensuring non-commutative layers do not entangle the separated components and maintaining scalability under hardware constraints.
In this work, we show that such a separation is not only possible but offers clear scaling advantages. We introduce sVQNHE, a variational framework in which a neural network (NN) models the amplitude distribution and a shallow diagonal quantum circuit learns the phase (sign) structure through neural-guided layer-wise optimization. To enable stacking multiple phase layers without collapse, since diagonal two-qubit gates alone support only quadratic interactions, we insert shallow, classically simulatable non-diagonal gates (e.g., a layer of $R_{y}$ gates) via a neural-guided amplitude transfer: the NN’s learned distribution is iteratively approximated onto these gates through classical simulations, preserving the decomposition by isolating amplitude uploads from phase encoding. This targeted quantum phase learning ensures that (i) dramatically reduced measurement cost due to commuting Pauli terms in the circuit, (ii) a well-conditioned optimization landscape that mitigates barren plateaus, (iii) significantly improved noise resilience from controlled circuit depth and NN mitigation, and (iv) more robust shot-based estimation of observables compared to the other non-unitary hybrid quantum-neural method. Notably, this design does not diminish quantum expressivity—neural amplitude modeling and quantum phase learning are jointly trained to reproduce nontrivial many-body states and combinatorial structure, representing a key innovation that bridges classical deep learning’s function approximation power with quantum hardware’s interference and entangling capabilities.
We demonstrate sVQNHE’s effectiveness across strongly correlated quantum many-body systems and large-scale combinatorial optimization. On the 6-qubit frustrated J1-J2 model, sVQNHE reduces mean absolute energy error by $98.9\%$ and variance by $99.6\%$ versus the classical NN baselines, converging $\sim 19\times$ faster than hardware-efficient VQE. It also resolves molecular ground states such as the $H_{2}O$ molecule and demonstrates robustness under depolarizing noise and finite sampling. In the Maximum Clique problem (135 vertices, p=0.5), sVQNHE using only 10 qubits via Pauli-correlation encoding, substantially outperforms greedy heuristics, and a recently proposed Free-Energy Machine [shen2025free] (FEM), thereby confirming the non-trivial contribution of its quantum phase-learning module in highly constrained, rugged landscapes. Furthermore, on a MaxCut instance with 1,485 vertices, it achieves the same performance as the state-of-the-art, classical FEM using only 12 qubits. These results highlight sVQNHE’s enhanced generalization across many-body physics and diverse combinatorial optimization challenges, including problems where classical solvers like FEM face limitations due to mean-field approximations. This provides evidence that explicit quantum phase learning, tightly coupled with classical amplitude modeling, constitutes a principled paradigm for scalable hybrid algorithms. By optimizing the co-learning strategy on the amplitude and phase at the representation level, sVQNHE offers a pathway toward reliable near-term quantum utility in quantum chemistry, materials science, and optimization, enabling NISQ-era devices to deliver meaningful performance before the arrival of full fault tolerance.
@@ Results
Figure 1: (a) The algorithm seamlessly integrates quantum circuits with neural networks. The quantum circuits interleaves diagonal layers ($W_{1}\cdots W_{n}$) with shallow, simultable layers ($G_{1}\cdots G_{n}$). Parameters of the $G$ layers are optimized classically using a gradual transfer mechanism, while those of the $W$ layers are adjusted via the parameter shift rule on the actual hardware. Leveraging the commuting properties of the diagonal layers, the number of measurements needed to compute gradients for the $d$ parameters of $W$ drops from $\mathcal{O}(2d)$ to $\mathcal{O}(1)$. (b) Illustration of sVQNHE’s advantages: The green dashed circle represents the coverage of the hybrid quantum-classical scheme. The star represents the optimal solution. (b1) describes the suboptimal situation, that is, the PQC cannot connect the solution region with the optimal solution; (b2) describes the potential quantum advantage. Similar to (b1), but the star is now inside the solution space, indicating successful coverage. (b3) Adjustable ansatz shows a transition from (b1) to (b2) via parameter adjustment, including gradient shading to indicate the evolution. (c) Illustration of the energy optimization process using the hierarchical optimization scheme. When entering the next layer, the local optimal solution will jump out due to the newly introduced parameters.
We introduce sVQNHE, a hybrid quantum-classical variational algorithm designed to address key limitations of conventional VQAs, including high measurement cost, barren plateaus, limited expressive power with shallow circuits, and noise-induced errors. This framework builds on the original VQNHE approach by introducing a layer-wise, phase-amplitude decoupling strategy[westerhout2023many] (see Figure 1) that enhances trainability, scalability, and sampling robustness.
@@
[…]
where $F_{L}$ denotes a post-processing operator defined by a classical NN. The following paragraphs explain the design of the hybrid ansatz and the associated bidirectional optimization strategy. Finally, the algorithm is recapped in Alg. 1.
[…]
Classical Neural Network for Amplitude Modeling. The classical neural network is constrained to produce non-negative outputs, and is used to model the amplitude $|F=\sum_{s}f(s)\ket{s}\bra{s}$ of the quantum state in the computational basis. With the phase information delegated to the quantum circuit, the NN is focused on learning the probability distribution. Previous works have shown that classical neural networks are highly effective in this task, particularly for systems with smooth or structured amplitude landscapes [westerhout2020generalization, westerhout2023many, szabo2020neural], relevant for many physical applications.
Bidirectional Feedback Loop: Decoupling Amplitude and Phase. At each iteration $i$, a new block of parametrized quantum layers $(G_{i},W_{i})$ is added to the circuit. The process follows two key steps. (1) Forward Initialization Step: The classical NN operator $F$ from the previous iteration is approximated by the shallow quantum layer $G_{i}$, via a tractable classical optimization procedure as detailed in the Method section. This ensures that $G_{i}$ inherits the amplitude structure previously learned by the NN. (2) Hybrid Optimization: Once $G_{i}$ has learn an approximation of $F$ and is loaded onto the quantum circuit, the neural operator $F$ is reset and jointly optimized with the parameters of the new diagonal gate layer $W_{i}$ by minimizing a loss function, such as the energy of a given Hamiltonian. This hybrid optimization simultaneously refines both the amplitude ($F$) and sign structure ($W_{i}$) in a tightly coupled quantum–classical feedback loop.
@@
6:Define diagonal operator $|F_{1}\leftarrow\sum_{s}f(s)\ket{s}\bra{s}$ $\triangleright$ $f(s)$: NN-predicted amplitude
7:Set $G_{1}\leftarrow H^{\otimes n}$ $\triangleright$ Uniform superposition
@@
9:while $\neg\text{converged}$ do $\triangleright$ Global convergence
@@
12:    Update $\theta^{g_{l}}$ to minimize $\|G_{l}-F_{l-1}\|$ $\triangleright$ Gradual Transfer Mechanism, see Method for detail
@@
14:   $E_{l}\leftarrow 0$ $\triangleright$ Reset layer energy
@@
[…]
@@
24:    Update parameters: $\theta^{w_{l}}$ $\triangleright$ Measurement Strategy, see Method for detail
25:   until $|\Delta E_{l}|<\epsilon_{\text{conv}}$ $\triangleright$ Energy convergence
@@
29:return $E_{L}$ $\triangleright$ Final optimized energy
Efficient Allocation of Quantum Resources
Traditional variational algorithms, such as the Variational Quantum Eigensolver (VQE) and the Quantum Approximate Optimization Algorithm (QAOA), typically employ fixed circuit ansatzes derived from physical intuitions like Trotterized time evolution[aaronson2009need, mcclean2021low]. These static architectures can limit performance when the ansatz fails to capture the entanglement or correlations required for the target state, often trapping optimization in local minima. Adaptive strategies like adapt-VQE attempt to address this by iteratively selecting gates from a predefined pool, but they remain constrained by the expressiveness of the pool and incur high measurement overhead due to repeated gate evaluations[tang2021qubit, grant2019initialization].
In contrast, sVQNHE employs a neural-guided adaptive hybrid ansatz (as defined in Equation (1)) that achieves an alternative and (potentially) more efficient allocation of quantum resources. The ansatz alternates between shallow and non-diagonal $G_{i}$ (e.g., $R_{y}$ rotations) and diagonal multi-qubit phase-encoding layers $W_{i}$ (e.g., $R_{z}$ and $R_{zz}$). Crucially, circuit growth proceeds in a layer-wise manner under the guidance of the classical neural post-processing module $F$. At each stage of optimization, the neural network receives the current output distribution from the quantum circuit and provides feedback that informs whether additional layers are required and, if so, how they should be structured to most effectively refine the approximation. This amplitude-guided, progressive deepening ensures that quantum resources are deployed selectively and only where necessary to capture residual discrepancies in the target distribution, avoiding unnecessary depth or parameter overhead.
Complementing this adaptive construction is a measurement strategy that exploits the commuting structure within each layer: the diagonal $W_{i}$ layers are fully commuting with extremely low measurement costs, while the parameters in shallow, simulatable $G_{i}$ layers are optimized classically without any measurement costs. As a result, gradient estimation requires substantially fewer quantum shots compared with the standard VQE algorithm employing the hardware-efficient or other general, entangling ansatzes, thereby reducing the overall measurement burden and enhancing scalability on devices with limited shot capacity. A detailed discussion of the measurement-efficient gradient protocol is provided in the Measurement Reduction part.
Even with relatively simple non-diagonal layers based on $R_{y}$ rotations, the chosen generator set $\{ZD_{2j},Y_{i}\}$ spans a Lie algebra of dimension $\mathfrak{su}(2^{n-1})\oplus\mathfrak{su}(2^{n-1})$ (Theorem 1), enabling the approximation of highly entangled states using only shallow circuit depth. Furthermore, the incorporation of neural post-processing helps maintain polynomial gradient variance in the loss landscape, $\mathrm{Var}[\partial_{\theta_{k}}\mathcal{L}]\in\Omega(1/\mathrm{poly}(n))$, thereby suppressing the emergence of barren plateaus during adaptive growth (Theorem 2). Proposition 1 additionally informs the design of non-commuting layer sequences, ensuring that added depth contributes meaningfully to expressiveness.
Consequently, sVQNHE dynamically assembles shallow yet highly expressive circuits whose reachable Hilbert space expands in a targeted, data-driven fashion until the ground-state wavefunction or desired combinatorial solution distribution is adequately represented (illustrated schematically in Figure 1(b3)). Empirical results across frustrated spin models (such as J1–J2 chains) and large-scale combinatorial optimization instances (including MaxCut and Maximum Clique under qubit-efficient encodings) demonstrate that sVQNHE consistently achieves high-fidelity output distributions using fewer quantum layers and markedly lower measurement costs than fixed-depth variational baselines, non-adaptive hybrid methods, or purely classical neural approaches, even for distributions that are classically difficult to sample due to strong entanglement and complex sign structure [troyer2005computational].
In summary, by integrating neural-directed adaptive growth, measurement-frugal layer designs, and Lie-algebra-enhanced shallow-circuit expressiveness, sVQNHE reallocates quantum resources far more efficiently than traditional fixed or purely adaptive variational frameworks. This resource-aware construction positions the method as a robust candidate for realizing practical quantum utility on near-term noisy quantum devices, where both qubit count and measurement budget remain severely constrained.
Reducing Measurement Costs and Enhancing Trainability
@@
The sVQNHE framework addresses these limitations through two synergistic innovations. First, it employs commuting diagonal gates, such as $R_{z}$ and $R_{zz}$ rotations, which enable simultaneous measurement of all relevant expectation values in one basis. This compresses the per-operator measurement cost to $\mathcal{O}(1)$, reducing the total overhead to $\mathcal{O}(\text{poly}(n))$ for $\text{poly}(n)$ operators (see details in the method section). This stands in stark contrast to the multiplicative $\mathcal{O}(2d\cdot\text{poly}(n))$ scaling of traditional VQNHE and conventional VQAs, allowing sVQNHE to tackle large-scale problems efficiently.
Secondly, our framework incorporates layer-wise neural guidance to decompose the optimization process into manageable subproblems, addressing a critical limitation of traditional layer-wise VQAs[liu2022layer]. In conventional layer-wise VQAs, the absence of a global perspective on the optimization landscape renders sequential layer optimization prone to converging at local minima. In contrast, the sVQNHE framework employs a classical neural network to guide the optimization of each quantum layer, producing a hybrid state $|\psi_{l,f}\rangle$ that is more expressive than the state $|\psi_{l}\rangle$ obtained from hierarchical quantum circuits alone. This neural guidance not only helps avoid suboptimal local minima but also focuses parameter updates on one specific block each time, such as $W_{l}$ at the $l$-th iteration. This approach, supported by the benign landscape of Theorem 2, restricts the effective parameter space per iteration (e.g., block $W_{l}$), preventing convergence to local minima and maintaining scalability. Together, these strategies circumvent the dual bottlenecks of measurement overhead and barren plateaus, establishing a scalable pathway for variational quantum algorithms in practical applications.
@@
[…]